OpenStudy (anonymous):

Is this following sequence bounded? And does it converge? If so find the limit: Xn=( n^2cos(n*pi))/(2n^2 + n)

OpenStudy (anonymous):

\[x _{n}= (n ^{2} \cos(n \Pi))\div(2n ^{2} + n)\]

OpenStudy (kainui):

cos(n*pi) = (-1)^n for integers, so might be helpful, I don't know.

OpenStudy (anonymous):

I'm not sure what to do thats why im a little confused

OpenStudy (kainui):

I honestly never remembered the tests for these, I would sometimes just try to look at it in different ways and be able to more or less figure it out through reckoning.

OpenStudy (kainui):

Like, obviously the top is oscillating between + and - and the limit of a single term will approach 1/2. Seems like it would be divergent.

OpenStudy (kainui):

But maybe I'm missing something, it's kind of tricky looking. Good luck. Try looking at Paul's online notes for calculus 2.

OpenStudy (anonymous):

It's a practice for an upcoming test .. Thankss anyways I'll check Paul' online notes

OpenStudy (kainui):

Hmmm... Wolfram Alpha says the root and ratio test are inconclusive. http://www.wolframalpha.com/input/?i=infinite%20sum%20(%20n%5E2cos(n*pi))%2F(2n%5E2%20%2B%20n)%20&t=crmtb01 Just an extra tool to help you practice for your test.

OpenStudy (anonymous):

So its means that the sequence isnt bounded ?

OpenStudy (kainui):

Yeah. As you approach infinity you will either get a -1 or +1 times your term because of the cosine term, right?

OpenStudy (kainui):

So remove that cosine part and just take the limit as it approaches infinity and the divergence test will show you that if it was just a positive (or negative) term it would fail, since the limit will be 1/2. So if you're "at" infinity and have 1/2 with every other number oscillating between +1 and -1 then eventually you'll just be at: +1/2-1/2+1/2-1/2+1/2-1/2+... which doesn't converge for the same reason that \[\lim_{x \rightarrow \infty} \cos(x)\] doesn't converge.

OpenStudy (kainui):

In fact, it'll look almost exactly like a more discrete form of this limit. \[\lim_{x \rightarrow \infty} \frac{ 1 }{ 2} \cos(x)\]

OpenStudy (anonymous):

I think I'm starting to get it .. So that means that the fucntion will oscillate between -.5 and . 5 AND DIVERGE TO INFINITY ? Sorry for the caps loll

OpenStudy (kainui):

Well, not necessarily diverge TO infinity but doesn't converge AT infinity because at infinity you don't know if it's at +1 or -1, but it's not infinity itself.

OpenStudy (anonymous):

I meant when n goes to infinity it oscilllates between -1and 1 but never converges to a limit

OpenStudy (kainui):

Yeah.

OpenStudy (anonymous):

Cool Thanks :)